# Supercell Wannier functions and faithful low-energy model for Bernal bilayer graphene

### v1

@@ Introduction.
In the present manuscript, we construct a fully symmetric lattice model that captures the low-energy physics of BBG and circumvents above obstacles. We reformulate the (interacting) physics in BBG and related rhombohedral graphene stacks at low electronic densities in terms of supercell Wannier functions (SWFs) that are localized on length scales of $L_{s}=5...50\,\mathrm{nm}$, which yields an elegant real-space analogue of the full momentum-space formalism adopted in the continuum limit, where the locality and the effective length scales of emergent phenomena are hidden. We demonstrate that spectral features and the local Berry curvature of the of the valley-local flat bands near $K^{\nu}$ can be faithfully represented by two $p_{\pm}$-orbitals per valley that span the $U(4)\otimes U(4)$ symmetry space of spin, valley and layer degrees of freedom. The SWFs therefore connect the atomic representation of the electronic structure as obtained from first-principles, e.g. via density functional theory (DFT) & Wannierization of the carbon $p_{z}$-orbitals [44], with a mesoscopic description of flavor-polarized quasiparticles that shape the emergent physics at low electronic densities. This allows to downfold realistic (dual-gated) Coulomb interactions to the effective Wannier basis including orbitally resolved short- and long-ranged interaction components [61, 62]. Our work not only demonstrates that the low-energy physics in BBG is governed by weak electron-electron interactions, but at the same time provides a natural real-space description that bridges to the emergent physics in twisted graphene multilayers.
@@
To circumvent a full momentum-space formalism of the low-energy physics in BBG and related small Fermi pocket systems, we propose a real-space supercell construction inspired by the moiré pattern in twisted (graphene) multilayers [63, 64, 65, 66], albeit without a twist between adjacent layers. Instead of describing the crystal structure of BBG in terms of its primitive unit cell with Bravais lattice vectors $\bm{a}_{1}=(0,1)^{\mathrm{T}}$ and $\bm{a}_{2}=(1/2,\sqrt{3}/2)^{\mathrm{T}}$, we set up a superlattice with rescaled lattice vectors $\bm{A}_{i}=n_{s}\bm{a}_{i}$  [^1] In principle, the construction works for all supercells that are the result of a linear integer map applied to the primitive lattice vectors, i.e., $(\bm{A}_{1},\bm{A}_{2})^{T}=\mathcal{N}\cdot(\bm{a}_{1},\bm{a}_{2})^{T}$, with $\mathcal{N}$ a $2\times 2$ matrix with integer entries., where $n_{s}$ denotes the linear scaling factor of the supercell. As shown in Fig. 1 (a), the superlattice contains $n_{s}^{2}$ primitive unit cells on the atomic scale and is characterized by the mesoscopic length scale $L_{s}=|\bm{A}_{i}|=n_{s}|\bm{a}_{i}|$. The associated Brillouin zone (BZ) is shrunk to a mini-BZ recovering momentum-space periodicity with respect to the reciprocal lattice vectors $\bm{G}_{i}=\bm{g}_{i}/n_{s}$ as indicated in Fig. 1 (b). The high-symmetry points of the mini-BZ are labeled according to their counterparts in the primitive BZ. Depending on the choice of the scaling factor $n_{s}$, the original graphene valleys $K^{\nu}$, $\nu=\pm$ are mapped to either the $\Gamma_{s}$-point [$\mathrm{mod}(n_{s},3)=0$] as shown in Fig. 1 (c), or onto the supercell $K_{s}^{\nu}$-points [$\mathrm{mod}(n_{s},3)\neq 0$].
The Hamiltonian of the supercell is constructed starting from a first-principle Hamiltonian defined in the primitive cell of BBG that is obtained from DFT & Wannierization [44]. The microscopic Wannier Hamiltonian is therefore written in the basis of the four carbon $p_{z}$-orbitals centered at the sublattices $X=(A_{1},B_{1},A_{2},B_{2})$, whereas the supercell Hamiltonian comprises $4n_{s}^{2}$ orbitals on the atomic scale, see Supplementary Material for technical details of the Hamiltonian setup. Its symmetries $\mathcal{G}^{0}$ are carried over from the microscopic model, i.e., $\mathcal{G}^{0}=C_{3v}\otimes SU(2)\otimes\mathcal{T}$. Here, $C_{3v}$ is the point group of the superlattice that contains a three-fold rotation around the $z$-axis and three vertical mirror planes, the $SU(2)$-symmetry acts in spin space and $\mathcal{T}$ denotes time-reversal symmetry. At low electronic densities, the system further possesses an approximate valley conservation $U_{\nu}(1)$ that is reflected in an orbital-dependent phase factor $e^{i\bm{K}^{\nu}\bm{r}}$ of the respective Bloch states. In the supercell picture, the valley character of the Bloch states $|\psi_{\bm{k}b}\rangle$ can be disentangled by evaluating expectation values $\langle\mathcal{V}\rangle=\langle\psi_{\bm{k}b}|\mathcal{V}|\psi_{\bm{k}b}\rangle$ of a Haldane-like operator $\mathcal{V}$ [68, 69], see Supplementary Material. As demonstrated in Fig. 1 (c), the back-folded bands in the mini-BZ of the supercell originating from valley $K^{\nu}$ indeed carry a well-defined expectation value $\langle\mathcal{V}\rangle\sim\nu$, whereas remote energy states originating from states near $\Gamma$ show no distinct valley polarization.
@@
The periodic supercell description of the atomic and electronic structure of BBG allows to construct a low-energy model for BBG from first-principles by finding a minimal Wannier representation that captures the spectral weight and Berry curvature of the valence and conduction bands near charge neutrality. To construct localized SWFs $\chi_{\bm{R}\zeta\nu}(\bm{r})$ in the presence of finite displacement fields, we apply the single-shot projection technique [70, 71, 72], see Supplementary Material for technical details. Choosing $\mathrm{mod}(n_{s},3)=0$ allows trial wave functions $g_{\bm{R}\zeta\nu}(\bm{r})$ to be trivially symmetric (i.e., without phase factors) under the model’s point group symmetries when sampled from Bloch states at $\Gamma_{s}$, i.e. $g_{\bm{R}\zeta\nu}(\bm{r})\sim\psi_{\Gamma_{s}\zeta\nu}(\bm{r})$. States originating from the valence (conduction) band manifold are polarized on the non-dimer orbitals $B_{2}$ $(A_{1})$ as labeled by $\zeta=\pm$. Due to the emergent valley conservation $U_{\nu}(1)$ and the presence of time-reversal symmetry $\mathcal{T}$ and spin-$SU(2)$ symmetry, each state at $\Gamma_{s}$ is fourfold degenerate and we further separate Bloch states according to their valley polarization $\nu=\pm$. The trial orbitals $g_{\bm{R}\zeta\nu}(\bm{r})$ transform as eigenstates of $C_{3v}$ with eigenvalue $w^{\zeta\nu}$, where $w=e^{i2\pi/3}$, and therefore resemble $p_{\pm}=p_{x}\pm ip_{y}$-orbitals with angular momentum $L_{z}=\nu\zeta$.
@@
In Fig. 2 (a) we show the valley-polarized SWFs $\chi_{\bm{R}\zeta+}(\bm{r})$ obtained in a BBG supercell with scaling factor $n_{s}=24$ and interlayer potential difference of $\Delta=60$ meV. However, we stress that the SWF construction presented explicitly for BBG can be generalized to arbitrary rhombohedral graphene stacks, see Supplementary Material for a discussion on monolayer and ABC trilayer graphene. The SWFs are decomposed into contributions from the microscopic sublattice- and layer degrees of freedom $X$: $\chi_{\bm{R}\zeta\nu}(\bm{r})=\{\chi^{X}_{\bm{R}\zeta\nu}(\bm{r})\}$. The amplitude of the SWF with quantum number $\zeta$ is enhanced on the non-dimer orbital $B_{2}$ $(A_{1})$ and decays exponentially on length scales associated to the BBG supercell lattice constant $L_{s}$. To resolve the relative phase windings of the SWFs on the individual sublattices, we separate contributions of the atomic valley phase factor from the supercell envelope function $\chi_{\zeta\nu}(\bm{r})=e^{i\bm{K}_{\nu}\bm{r}}W_{\zeta\nu}(\bm{r})$. The former features $C_{3v}$ eigenvalues of $(w^{-1},1,1,w)^{\nu}$ on the different sublattices, whereas the supercell envelope function $W_{\zeta\nu}(\bm{r})$ has eigenvalues $(w^{(\zeta+1)},w^{\zeta},w^{\zeta},w^{(\zeta-1)})^{\nu}$, see Supplementary Material. Therefore, the total SWFs have the same $C_{3v}$ eigenvalue $w^{\zeta\nu}$ consistent with the $p_{\pm}$-orbital character of the trial functions. In particular, the sublattice components of the SWFs each have a relative phase winding of $2\times 2\pi$, which is reminiscent of the wave functions of chiral quasiparticles $\psi_{\bm{k}\zeta\nu}(\bm{r})\sim e^{i\bm{k}\bm{r}}(1,\zeta e^{Ji\nu\varphi_{\bm{k}}})^{\mathrm{T}}$, where $\varphi_{\bm{k}}=\arg(k_{x}+ik_{y})$ [73, 45] and $J=2$. The latter are well-known to govern the low-energy physics in the continuum formulation of rhombohedral multilayer graphene [73, 45], where adiabatic propagation along a closed orbit yields a Berry phase of $J\pi$ for $J$-layered rhombohedral graphene. This manifests in a half-integer valued valley Chern number $C_{\nu}=J\zeta\nu/2$ [74, 75] characterized by an accumulation of Berry curvature $\Omega_{\nu}(\bm{k})$ in the vicinity of $K^{\nu}$. We demonstrate in Fig. 2 (d) that the SWFs accurately reproduce the Berry curvature of the original Bloch states near charge neutrality. Due to the momentum-space periodicity of the SWFs and the connectedness of the lowest valence (conduction) bands, the Berry curvature of the original Hamiltonian can only be captured in a constrained energy window. We find that the localized SWFs form a Haldane-like model within each valley featuring a valence (conduction) band valley Chern number $C_{\nu}=\pm 2\nu$ and thus represent an atomic limit in each valley. As time-reversal symmetry $\mathcal{T}$ connects different valley sectors $\chi_{\bm{R}\zeta\bar{\nu}}(\bm{r})=\chi^{*}_{\bm{R}\zeta\nu}(\bm{r})$, valence (conduction) bands of the full model retain Chern number $C=0$. In contrast to topologically (or chirally) obstructed low-energy models [76], the SWFs capture the entire spectral weight of the Bloch states near charge neutrality. The overlaps of the SWFs with the original Bloch states satisfy $\braket{\chi_{\bm{k}\zeta\nu}}{\psi_{\bm{k}b}}=1$ around $\Gamma_{s}$ as shown in Fig. 2 (c), while decreasing at the BZ boundary. Further evidence for the valley-polarized SWFs representing an atomic limit at low densities concerns the exponential decay of the hopping parameters $\mathsf{t}_{(\bm{R}\zeta\nu)(\bm{R}^{\prime}\zeta^{\prime}\nu^{\prime})}$ as function of distance, see the right panel of Fig. 2 (c). Due to the explicit symmetries $U_{\nu}(1)$, $\mathcal{T}$ of the low-energy model, the hopping terms fulfill $\mathsf{t}_{(\bm{R}\zeta\bar{\nu})(\bm{R}^{\prime}\zeta^{\prime}\bar{\nu})}=[\mathsf{t}_{(\bm{R}\zeta\nu)(\bm{R}^{\prime}\zeta^{\prime}\nu)}]^{*}$ and $\mathsf{t}_{(\bm{R}\zeta\nu)(\bm{R}^{\prime}\zeta^{\prime}\bar{\nu})}=0$, such that the intra- and inter-orbital hopping terms in a single valley determine the full model. In Fig. 2 (c), we further demonstrate that the energetics of the valence (conduction) bands around charge neutrality are accurately reproduced by the SWFs. As shown in Fig. 2 (b), the effect of an external displacement field can be modeled by interpolating hopping terms at finite interlayer potential differences, while reproducing all characteristics of the low-energy bands including trigonal warping and the size of the single-particle band gap.
@@
Figure 3: Strength of Coulomb interactions in the Wannier basis. (a) Density-density vertex element $U^{0R0R}$ as a function of distance $r$ for various gate distances (indicated as dash-dotted vertical lines of the respective color). The supercell size is chosen as $n_{s}=24$. (b) On-site Hubbard-Kanamori interaction components as a function of supercell size $n_{s}$. The left axis refers to Hubbard-$U$ terms and the right axis to Hund-$J$ couplings.
Besides its major advantage of a symmetric lattice description, formulating the low-energy physics of BBG in terms of SWFs has the benefit of obtaining electron-electron interactions from first principles through the intrinsic connection to the atomistic basis of carbon $p_{z}$-orbitals. We assume a realistic device geometry [6, 1], where a BBG sheet is placed inbetween hBN spacer layers of thickness $\xi$ that determine the dielectric environment $\epsilon\approx 4$ [23]. The hBN/BBG/hBN stack is enclosed with metallic (graphite) gates. These assumptions determine the overall strength $\alpha/\epsilon$ and characteristic distance $\xi$ of the long-range part of the Coulomb interaction. The short-range behavior of the Coulomb repulsion follows an Ohno form, with parameters for graphitic systems known from literature [61, 62]. This fixes the microscopic parametrization of the Coulomb interaction $V(r)$, see Supplementary Materials for details. The supercell Wannier construction allows to project this first-principle $V(r)$ to the SWF basis (see Fig. 3). First, we analyze density-density type interactions obtained by evaluating the Coulomb integral
@@
$J_{0000}^{\alpha_{1}\dots\alpha_{4}}=\!\iint\!\!\mathrm{d}\bm{r}\mathrm{d}\bm{r}^{\prime}\,\chi^{*}_{0\alpha_{1}}(\bm{r})\chi^{*}_{0\alpha_{2}}(\bm{r}^{\prime})\chi_{0\alpha_{4}}(\bm{r}^{\prime})\chi_{0\alpha_{3}}(\bm{r})V(|\bm{r}-\bm{r}^{\prime}|)\,.$ (2)
as well as their dependence on the supercell scaling factor $n_{s}$, see Fig. 3 (b). In the enlarged pseudospin space including spin, valley and orbital (layer) degrees of freedom, the presence of time-reversal symmetry mandates four unique types of local interactions: (i) intra-orbital ($\zeta=\pm$ for $1,2$, respectively) $U^{(1,2)}=U^{(\zeta\nu)(\zeta\nu)(\zeta\nu)(\zeta\nu)}$ and inter-orbital $U^{(3)}=U^{(\zeta\nu)(\bar{\zeta}\nu)(\zeta\nu)(\bar{\zeta}\nu)}$ Hubbard interactions, (ii) inter-orbital, intra-valley coupling $J^{(1)}=J^{(\zeta\nu)(\bar{\zeta}\nu)(\bar{\zeta}\nu)(\zeta\nu)}$, (iii) intra-orbital, inter-valley coupling $J^{(2)}=J^{(\zeta\nu)(\zeta\bar{\nu})(\zeta\bar{\nu})(\zeta\nu)}$ and (iv) inter-orbital and inter-valley coupling $J^{(3)}=J^{(\zeta\nu)(\bar{\zeta}\bar{\nu})(\bar{\zeta}\bar{\nu})(\zeta\nu)}$. The local Hubbard interaction terms scale like $U^{(i)}\sim 1/n_{s}$ as expected from the Gaussian shape of the SWFs [77]. In particular, we find that for all supercell sizes $n_{s}=15\dots 30$ considered within the scope of this work, the minimal model of BBG remains in the weakly-coupled regime $\max_{i}(U^{(i)})<W$, where $W$ denotes the supercell bandwidth. For the supercell scaling factor of $n_{s}=24$ used throughout the manuscript, we estimate the ratio of local interaction strength over bandwidth as $U/W\approx 0.15$. Hund’s coupling terms are suppressed by a factor of $\sim\!\!10$ compared to the intra-orbital and inter-orbital Hubbard interactions. The dominant Hund’s interaction term in the SWF basis is the intra-orbital, inter-valley coupling $J^{(2)}$.

### v2

# Supercell Wannier functions and a faithful low-energy model for Bernal bilayer graphene

@@ Introduction.
In the present manuscript, we construct a fully symmetric lattice model that captures the low-energy physics of BBG and circumvents above obstacles. We reformulate the (interacting) physics of BBG and related rhombohedral graphene stacks at low electronic densities in terms of supercell Wannier functions (SWFs) that are localized on intermediate length scales $a_{0}<L_{s}<1/|\bm{q}_{\mathrm{FS}}|\sim 60a_{0}$ (with $|\bm{q}_{\mathrm{FS}}|$ the characteristic extent of the Fermi pockets). By comprising several hundreds of carbon atoms on the atomic carbon-carbon scale $a_{0}$, the SWFs represent a coarse-grained description of rhombohedral graphene stacks that links the atomic representation of the electronic structure as obtained from first-principles, e.g. via density functional theory (DFT) & Wannierization of the carbon $p_{z}$-orbitals [44] with a mesoscopic description at intermediate length scales $L_{s}$. Therefore, the SWFs yield an elegant real-space analogue of the full momentum-space formalism adopted in the continuum limit, where the locality and the effective length scales of emergent phenomena are hidden. We demonstrate that spectral features and the local Berry curvature of the valley-local flat bands near $K^{\nu}$ can be faithfully represented by two $p_{\pm}$-orbitals per valley that span the $U(4)\otimes U(4)$ symmetry space of spin, valley and layer degrees of freedom. This allows to downfold realistic (dual-gated) Coulomb interactions to the effective Wannier basis including orbitally resolved short- and long-ranged interaction components [61, 62]. Our work not only demonstrates that the low-energy physics in BBG is governed by weak electron-electron interactions, but at the same time provides a natural real-space description that bridges to the emergent physics in twisted graphene multilayers.
@@
To circumvent a full momentum-space formalism of the low-energy physics in BBG and related small Fermi pocket systems, we propose a real-space supercell construction inspired by the moiré pattern in twisted (graphene) multilayers [63, 64, 65, 66], albeit without a twist between adjacent layers. For the former, Wannier models have become state of the art for studies going beyond the single-particle picture [67, 68, 69, 70, 71, 72, 73]. Instead of describing the crystal structure of BBG in terms of its primitive unit cell with Bravais lattice vectors $\bm{a}_{1}=(\sqrt{3}/2,1/2)^{\mathrm{T}}$ and $\bm{a}_{2}=(0,1)^{\mathrm{T}}$, we set up a superlattice with rescaled lattice vectors $\bm{A}_{i}=n_{s}\bm{a}_{i}$  [^1] In principle, the construction works for all supercells that are the result of a linear integer map applied to the primitive lattice vectors, i.e., $(\bm{A}_{1},\bm{A}_{2})^{T}=\mathcal{N}\cdot(\bm{a}_{1},\bm{a}_{2})^{T}$, with $\mathcal{N}$ a $2\times 2$ matrix with integer entries., where $n_{s}$ denotes the linear scaling factor of the supercell. As shown in Fig. 1 (a), the superlattice contains $n_{s}^{2}$ primitive unit cells on the atomic scale and is characterized by the mesoscopic length scale $L_{s}=|\bm{A}_{i}|=n_{s}|\bm{a}_{i}|$. The associated Brillouin zone (BZ) is shrunk to a mini-BZ recovering momentum-space periodicity with respect to the reciprocal lattice vectors $\bm{G}_{i}=\bm{g}_{i}/n_{s}$ as indicated in Fig. 1 (b). The high-symmetry points of the mini-BZ are labeled according to their counterparts in the primitive BZ. Depending on the choice of the scaling factor $n_{s}$, the original graphene valleys $K^{\nu}$, $\nu=\pm$ are mapped to either the $\Gamma_{s}$-point [$\mathrm{mod}(n_{s},3)=0$] as shown in Fig. 1 (b), or onto the supercell $K_{s}^{\nu}$-points [$\mathrm{mod}(n_{s},3)\neq 0$]. To construct the Hamiltonian of the supercell from first-principles, we first consider an ab-initio Hamiltonian of BBG formulated in the basis of the four carbon $p_{z}$-orbitals centered at the sublattices $X=(A_{1},B_{1},A_{2},B_{2})$ as routinely derived from DFT & Wannierization [44]. As the supercell construction itself preserves translational invariance on the carbon-carbon bond scale, the Hamiltonian associated to the primitive cell can be mapped to the supercell geometry. Hence we avoid costly DFT simulations of all $4n_{s}^{2}$ carbon $p_{z}$-orbitals that reside within the supercell, see Supplementary Material (SM) [75] for technical details. The symmetries of the supercell Hamiltonian $\mathcal{G}^{0}$ are carried over from the microscopic model, i.e., $\mathcal{G}^{0}=C_{3v}\otimes SU(2)\otimes\mathcal{T}$. Here, $C_{3v}$ is the point group of the superlattice that contains a three-fold rotation around the $z$-axis and three vertical mirror planes, the $SU(2)$-symmetry acts in spin space and $\mathcal{T}$ denotes time-reversal symmetry. At low electronic densities, the system further possesses an approximate valley conservation $U_{\nu}(1)$ that is reflected in an orbital-dependent phase factor $e^{i\bm{K}^{\nu}\bm{r}}$ of the respective Bloch states. In the supercell picture, the valley character of the Bloch states $|\psi_{\bm{k}b}\rangle$ can be disentangled by evaluating expectation values $\langle\mathcal{V}\rangle=\langle\psi_{\bm{k}b}|\mathcal{V}|\psi_{\bm{k}b}\rangle$ of a Haldane-like operator $\mathcal{V}$ [76, 77], see SM [75]. As demonstrated in Fig. 1 (c), the back-folded bands in the mini-BZ of the supercell originating from valley $K^{\nu}$ indeed carry a well-defined expectation value $\langle\mathcal{V}\rangle\sim\nu$, whereas remote energy states originating from states near $\Gamma$ show no distinct valley polarization.
@@
The periodic supercell description of the atomic and electronic structure of BBG allows to construct a low-energy model for BBG from first-principles by finding a minimal Wannier representation that captures the spectral weight and Berry curvature of the valence and conduction bands near charge neutrality. In this work, we focus on the intermediate regime $L_{s}<1/|\bm{q}_{\mathrm{FS}}|$, where $|\bm{q}_{\mathrm{FS}}|$ denotes the (inverse) linear extent of the Fermi surface pockets around $K^{\nu}$. In this regime, only one band per valley contributes to the Fermi surface close to the VHSs and the artificial length scale set by the supercell does hence not infer with emergent length scales associated to possible electronic ordering, see Fig. 1 in the SM [75]. To construct localized SWFs $\chi_{\bm{R}\zeta\nu}(\bm{r})$ in the presence of finite displacement fields, we apply the single-shot projection technique [78, 79, 68], see SM [75] for technical details. Choosing $\mathrm{mod}(n_{s},3)=0$ allows trial wave functions $g_{\bm{R}\zeta\nu}(\bm{r})$ to be trivially symmetric (i.e., without phase factors) under the model’s point group symmetries when sampled from Bloch states at $\Gamma_{s}$, i.e. $g_{\bm{R}\zeta\nu}(\bm{r})\sim\psi_{\Gamma_{s}\zeta\nu}(\bm{r})$. States originating from the valence (conduction) band manifold are polarized on the non-dimer orbitals $B_{2}$ $(A_{1})$ as labeled by $\zeta=\pm$. Due to the emergent valley conservation $U_{\nu}(1)$ and the presence of time-reversal symmetry $\mathcal{T}$ and spin-$SU(2)$ symmetry, each state at $\Gamma_{s}$ is fourfold degenerate and we further separate Bloch states according to their valley polarization $\nu=\pm$. The trial orbitals $g_{\bm{R}\zeta\nu}(\bm{r})$ transform as eigenstates of $C_{3v}$ with eigenvalue $w^{\zeta\nu}$, where $w=e^{i2\pi/3}$, and therefore resemble $p_{\pm}=p_{x}\pm ip_{y}$-orbitals with angular momentum $L_{z}=\nu\zeta$.
@@
In Fig. 2 (a) we show the valley-polarized SWFs $\chi_{\bm{R}\zeta+}(\bm{r})$ obtained in a BBG supercell with scaling factor $n_{s}=24$ and interlayer potential difference of $\Delta=60$ meV. However, we stress that the SWF construction presented explicitly for BBG can be generalized to arbitrary rhombohedral graphene stacks, see SM [75] for a discussion on monolayer and ABC trilayer graphene. The SWFs are decomposed into contributions from the microscopic sublattice- and layer degrees of freedom $X$: $\chi_{\bm{R}\zeta\nu}(\bm{r})=\{\chi^{X}_{\bm{R}\zeta\nu}(\bm{r})\}$. The amplitude of the SWF with quantum number $\zeta$ is enhanced on the non-dimer orbital $B_{2}$ $(A_{1})$ and decays exponentially on length scales associated to the BBG supercell lattice constant $L_{s}$. To resolve the relative phase windings of the SWFs on the individual sublattices, we separate contributions of the atomic valley phase factor from the supercell envelope function $\chi_{\zeta\nu}(\bm{r})=e^{i\bm{K}_{\nu}\bm{r}}W_{\zeta\nu}(\bm{r})$. The former features $C_{3v}$ eigenvalues of $(w^{-1},1,1,w)^{\nu}$ on the different sublattices, whereas the supercell envelope function $W_{\zeta\nu}(\bm{r})$ has eigenvalues $(w^{(\zeta+1)},w^{\zeta},w^{\zeta},w^{(\zeta-1)})^{\nu}$, see SM [75]. Therefore, the total SWFs have the same $C_{3v}$ eigenvalue $w^{\zeta\nu}$ consistent with the $p_{\pm}$-orbital character of the trial functions. In particular, the sublattice components of the SWFs each have a relative phase winding of $2\times 2\pi$, which is reminiscent of the wave functions of chiral quasiparticles $\psi_{\bm{k}\zeta\nu}(\bm{r})\sim e^{i\bm{k}\bm{r}}(1,\zeta e^{Ji\nu\varphi_{\bm{k}}})^{\mathrm{T}}$, where $\varphi_{\bm{k}}=\arg(k_{x}+ik_{y})$ [80, 45] and $J=2$. The latter are well-known to govern the low-energy physics in the continuum formulation of rhombohedral multilayer graphene [80, 45], where adiabatic propagation along a closed orbit yields a Berry phase of $J\pi$ for $J$-layered rhombohedral graphene. This manifests in a half-integer valued valley Chern number $C_{\nu}=J\zeta\nu/2$ [81, 82] characterized by an accumulation of Berry curvature $\Omega_{\nu}(\bm{k})$ in the vicinity of $K^{\nu}$. We demonstrate in Fig. 2 (d) that the SWFs accurately reproduce the Berry curvature of the original Bloch states near charge neutrality. Due to the momentum-space periodicity of the SWFs and the connectedness of the lowest valence (conduction) bands, the Berry curvature of the original Hamiltonian can only be captured in a constrained energy window. We find that the localized SWFs form a Haldane-like model within each valley featuring a valence (conduction) band valley Chern number $C_{\nu}=\pm 2\nu$ and thus represent an atomic limit in each valley. As time-reversal symmetry $\mathcal{T}$ connects different valley sectors $\chi_{\bm{R}\zeta\bar{\nu}}(\bm{r})=\chi^{*}_{\bm{R}\zeta\nu}(\bm{r})$, valence (conduction) bands of the full model retain Chern number $C=0$. In contrast to topologically (or chirally) obstructed low-energy models [83], the SWFs capture the entire spectral weight of the Bloch states near charge neutrality. The overlaps of the SWFs with the original Bloch states satisfy $\braket{\chi_{\bm{k}\zeta\nu}}{\psi_{\bm{k}b}}=1$ around $\Gamma_{s}$ as shown in Fig. 2 (c), while decreasing at the BZ boundary. Further evidence for the valley-polarized SWFs representing an atomic limit at low densities concerns the exponential decay of the hopping parameters $\mathsf{t}_{(\bm{R}\zeta\nu)(\bm{R}^{\prime}\zeta^{\prime}\nu^{\prime})}$ as function of distance, see the right panel of Fig. 2 (c). Due to the explicit symmetries $U_{\nu}(1)$, $\mathcal{T}$ of the low-energy model, the hopping terms fulfill $\mathsf{t}_{(\bm{R}\zeta\bar{\nu})(\bm{R}^{\prime}\zeta^{\prime}\bar{\nu})}=[\mathsf{t}_{(\bm{R}\zeta\nu)(\bm{R}^{\prime}\zeta^{\prime}\nu)}]^{*}$ and $\mathsf{t}_{(\bm{R}\zeta\nu)(\bm{R}^{\prime}\zeta^{\prime}\bar{\nu})}=0$, such that the intra- and inter-orbital hopping terms in a single valley determine the full model. In Fig. 2 (c), we further demonstrate that the energetics of the valence (conduction) bands around charge neutrality are accurately reproduced by the SWFs. As shown in Fig. 2 (b), the effect of an external displacement field can be modeled by interpolating hopping terms at finite interlayer potential differences, while reproducing all characteristics of the low-energy bands including trigonal warping and the size of the single-particle band gap.
@@
Figure 3: Strength of Coulomb interactions in the Wannier basis. (a) Density-density vertex element $U^{0R0R}$ as a function of distance $r$ for various gate distances (indicated as dash-dotted vertical lines of the respective color). The supercell size is chosen as $n_{s}=24$. (b) On-site Hubbard-Kanamori interaction components as a function of supercell size $n_{s}$. The left axis refers to Hubbard-$U$ terms and the right axis to Hund-$J$ couplings. The grey dashed line indicates the evolution of the bandwidth $W$ of the effective model.
Besides its major advantage of a symmetric lattice description, formulating the low-energy physics of BBG in terms of SWFs has the benefit of obtaining electron-electron interactions from first principles through the intrinsic connection to the atomistic basis of carbon $p_{z}$-orbitals. We assume a realistic device geometry [6, 1], where a BBG sheet is placed inbetween hBN spacer layers of thickness $\xi$ that determine the dielectric environment $\epsilon\approx 4$ [23]. The hBN/BBG/hBN stack is enclosed with metallic (graphite) gates. These assumptions determine the overall strength $\alpha/\epsilon$ and characteristic distance $\xi$ of the long-range part of the Coulomb interaction. The short-range behavior of the Coulomb repulsion follows an Ohno form, with parameters for graphitic systems known from literature [61, 62]. This fixes the microscopic parametrization of the Coulomb interaction $V(r)$, see SM [75] for details. The supercell Wannier construction allows to project this first-principle $V(r)$ to the SWF basis (see Fig. 3). First, we analyze density-density type interactions obtained by evaluating the Coulomb integral
@@
[…]
as well as their dependence on the supercell scaling factor $n_{s}$, see Fig. 3 (b). In the enlarged pseudospin space including spin, valley and orbital (layer) degrees of freedom, the presence of time-reversal symmetry mandates four unique types of local interactions: (i) intra-orbital ($\zeta=\pm$ for $1,2$, respectively) $U^{(1,2)}=U^{(\zeta\nu)(\zeta\nu)(\zeta\nu)(\zeta\nu)}$ and inter-orbital $U^{(3)}=U^{(\zeta\nu)(\bar{\zeta}\nu)(\zeta\nu)(\bar{\zeta}\nu)}$ Hubbard interactions, (ii) inter-orbital, intra-valley coupling $J^{(1)}=J^{(\zeta\nu)(\bar{\zeta}\nu)(\bar{\zeta}\nu)(\zeta\nu)}$, (iii) intra-orbital, inter-valley coupling $J^{(2)}=J^{(\zeta\nu)(\zeta\bar{\nu})(\zeta\bar{\nu})(\zeta\nu)}$ and (iv) inter-orbital and inter-valley coupling $J^{(3)}=J^{(\zeta\nu)(\bar{\zeta}\bar{\nu})(\bar{\zeta}\bar{\nu})(\zeta\nu)}$. The local Hubbard interaction terms scale like $U^{(i)}\sim 1/n_{s}$ as expected from the Gaussian shape of the SWFs [69]. In particular, we find that for all supercell sizes $n_{s}=15\dots 30$ considered within the scope of this work, the minimal model of BBG remains in the weakly-coupled regime $\max_{i}(U^{(i)})<W$, where $W$ denotes the supercell bandwidth. For the supercell scaling factor of $n_{s}=24$ used throughout the manuscript, we estimate the ratio of local interaction strength over bandwidth as $U/W\approx 0.15$. For $W\gg\Delta$, the quadratic nature of the bands leads to the ratio of Hubbard interaction over band width scaling like $U/W\sim n_{s}$. The scaling behavior of $U/W$ changes for $W\gtrsim\Delta$, because the band width is not solely determined by the supercell scaling factor $n_{s}$ in this case. Hund’s coupling terms are suppressed by a factor of $\sim\!\!10$ compared to the intra-orbital and inter-orbital Hubbard interactions. The dominant Hund’s interaction term in the SWF basis is the intra-orbital, inter-valley coupling $J^{(2)}$.